# An Analytical and Experimental Study of Distributed Uplink Beamforming in the Presence of Carrier Frequency Offsets

### v1

@@ IV-B Monte Carlo Simulation of SINR Performance
We simulate an OFDM transmission and reception chain based on the 802.11 standard [38] OFDM format. Out of $N_{sc}=64$ subcarriers, we use $48$ subcarriers for data transmission, $4$ for pilot, and the rest as null subcarriers. The model simulates $M=4$ dAPs with $N=16$ antennas per dAP, serving $K=4$ single-antenna users at the same time-frequency resource. Users transmit their signals in uplink with the same power $p_{\mathrm{ul}}$, where the ratio of $p_{\mathrm{ul}}$ to the noise variance $\sigma_{z}^{2}$ in the environment is $20$ dB. Users first send orthogonal uplink pilots. These pilots are composed of two consecutive long training sequence (LTS) symbols based on 802.11 format and will be used for channel estimation and combining weight calculation. The cyclic prefix length is $L_{cp}=16$, and QPSK modulation will be used, which is more robust to CFO compared to higher-order modulations. CFO is induced in the system in the form of realizations of the normal distribution as $\epsilon\sim\mathcal{N}(0,\beta^{2})$. After receiving signals and estimating channels, CBF is applied, and SINR and EVM are measured for each scenario.
@@
These simulations are conducted to compare the measured SINR—obtained by introducing CFO as random realizations—with the analytical SINR expression in (III-B2), where the CFO effects are modeled using the closed-form approximations in (17)–(22). A comparison is made in Fig. 3a, where both the measured SINR based on the Monte Carlo simulation (labeled as Sim) and the analytical SINR based on (III-B2) and (17)–(22) (labeled as Anlt) are shown for CBF. The CFO distribution is normal, with STD $\beta$ ranging from $0.5\%$ to $5\%$. For instance, when $\beta=2\%$, according to the characteristics of the normal distribution, over $99.7\%$ of the normalized CFO ($\epsilon$) values are within $3\beta=6\%$ of the mean, which means $\epsilon\in[-6\%,\,6\%]$. For a subcarrier spacing of $\Delta f=30$ kHz, $6\%$ normalized CFO means the offset of $\delta_{f}=1.8$ kHz from the true carrier frequency. This frequency offset is equivalent to $600$ parts per billion (PPB) with a carrier frequency of $2.4$ GHz in the unlicensed band or $500$ PPB with a carrier frequency of $3.6$ GHz in the CBRS band. This range of CFO can significantly affect the performance of the system. The results in Fig. 3a illustrate that SINR drops with increasing $\beta$, and this drop is faster for larger OFDM symbol indices. This is because the phase rotation caused by CFO linearly scales with the symbol index $n$ within an uplink slot with multiple OFDM symbols. It also demonstrates that the SINR expression (III-B2) with the CFO model (17)–(22) gives an appropriate estimate of the average SINR for different $\beta$ and $n$ values.
Although we assess the SINR in our analytical framework, EVM is often used as a performance metric to evaluate empirical datasets [39, 40]. SINR accounts for interference effects and serves to measure the channel quality indicator (CQI), which is utilized as the parameter for dynamically adjusting the modulation and coding scheme (MCS) [41, 42]. However, in a multi-user transmission scenario, measuring the power of IUI and separating it from the power of a desired signal and the power of noise is not feasible in most practical scenarios, subsequently making accurate SINR measurements difficult [43, 44]. Due to the practical challenges of SINR measurement, EVM can be a viable alternative metric in real-world experiments because of the feasibility of measurement. EVM can provide insight into the different sources of imperfections, including carrier leakage, IQ imbalance, non-linearity, and LO frequency errors [45, 46]. Moreover, in several wireless communication standards, the requirements on EVM are defined as part of the standard, such as IEEE 802.11a [47] and IEEE 802.16e [48], among others.
@@
Accordingly, we evaluate the root mean square error vector magnitude ($\mathrm{EVM}_{rms}$) in our Monte Carlo simulations, as shown in Fig. 3b, where $\mathrm{EVM}_{rms}$ is defined as in [45]:
@@ V Experiments
In this section, we first introduce the RENEW [31] indoor massive MIMO platform. We then describe the experimental setup and outline the datasets collected under various experimental conditions. Furthermore, we showcase several experimental results derived from these datasets along with comparisons with the analytical framework discussed in § III.
@@ V-A Hardware Setup
We use the RENEW programmable massive MIMO platform [31] to perform experiments. The RENEW base station uses the Iris SDR [34] as its building block, with each SDR supporting two RF chains. The base station includes a central hub that by default distributes clock and time trigger signals to up to 48 Iris SDRs (96 antennas) for time and frequency synchronization. At each Iris, both the RF and sampling clocks are derived from the clock distributed by the hub. Therefore, all antennas are locked at the carrier frequency $f_{c}$ in the form of a perfectly synchronized array. In this scenario, perfect time and frequency synchronization among antennas is achieved via the central hub. We denote this operational mode of the SDRs as HUB-mode. An Iris can also be used in a standalone mode to emulate a user. The RENEW platform is designed to operate at the CBRS spectrum ranging from 3.553.55-$3.7$ GHz.
Apart from receiving an external clock, each Iris is equipped with its own LO, a super-TCXO with a nominal $\pm 100$ PPB accuracy for the sub-$6$ GHz band. Through firmware settings, the Iris can be configured to use its own LO to generate the RF clock or to use the external clock from the central hub. To emulate a distributed antenna array with independent LOs, each Iris in the base station can be programmed to use its own LO, introducing a random CFO between the antennas. In this operation mode, the antennas still receive the time trigger signal from the central hub as their time reference. Time synchronization in distributed systems can be achieved through well-established protocols, such as IEEE 1588 Precision Time Protocol (PTP) or GPS-based methods [49, 50, 51, 52]. Distributing the time trigger across radios ensures time alignment, allowing us to isolate and focus solely on the impact of CFO on system performance. We denote the operation mode in which SDRs use their LO to derive the RF clock as LO-mode.
@@ V-B Experiment Design
We use the RENEWLab real-time software [53] to collect datasets in all of the configurations. The RENEWLab software runs on a host server that is connected to the RENEW BS through fiber. Through a configurable frame schedule, RENEWLab software orchestrates the transmission and reception of pilot and data OFDM symbols at users and from the BS antennas, respectively. Within each frame, one of the BS antennas transmits a synchronization beacon so that users can synchronize their frame time boundaries with the BS frame. We set up the frame schedule for the users so that each user sends time-orthogonal pilots and then all simultaneously transmit data OFDM symbols. We use the 802.11 LTS signal as the pilot symbol, which is used for channel estimation and beamforming matrix calculation. The format of the data OFDM symbols also follows the 802.11 standard [38], where the FFT size is $64$. Out of the $64$ subcarriers, $48$ subcarriers carry data, $4$ are pilot subcarriers, and the remaining are null subcarriers. The pilot subcarriers are used for residual phase offset correction. We try QPSK for modulating data subcarriers. The detailed parameters of the experiments are shown in Table I. RENEWLab framework enables the recording of pilot and data OFDM symbols received at BS antennas in a dataset file with HDF5 format [54]. The HDF5 datasets can be processed offline with the Python-based post-processing library in the RENEWLab software. Our data collection effort resulted in over 250 HDF5 files, with each file comprising 2000 frames, and each frame containing pilot symbols for each user followed by $10$ data OFDM symbols. Datasets are available on [32] for further research purposes.
@@ V-C2 Beamforming Performance Analysis
Across various multi-user configurations—where different user subsets were selected to capture the effects of high and low spatial correlation in beamforming—we measured the average estimated CFO between the dAPs. The results show that inter-dAP CFO values in these multi-user experiments ranged from $150$ to $350$ PPB, while intra-dAP synchronization remained consistently maintained, as expected. Notably, the maximum observed CFO across all multi-user measurements approached approximately $900$ PPB, which corresponds to a normalized CFO of $\epsilon=4.15\,\%$ when considering a subcarrier spacing of $\Delta f=78.125$ kHz. These results are consistent with our assumption of negligible ICI impact, given that normalized inter-dAP CFO variations remain within the $5\%$ range.
@@
For two multi-user configurations of $64\times 2$ and $64\times 4$ with distributed placement of APs, we processed the received data samples in both the HUB-mode and LO-mode, employing ZFBF to mitigate inter-user interference and recover the uplink data from each user. In the corresponding Monte Carlo simulations, we used the estimated CSI and CFO values obtained from the experimental measurements. The average per-user EVM-SNR measurements for these multi-user scenarios are presented in Fig. 9, comparing the performance of Dist-HUB versus Dist-LO scenarios. As illustrated in Fig. 9, the performance drop between the Dist-HUB and Dist-LO scenarios is notable. In the two-user configuration, there is a reduction of approximately $5$ dB in the average EVM-SNR, while this degradation escalates to approximately $9$ dB in the four-user setup. This performance gap highlights the increased sensitivity of higher-order multi-user MIMO configurations to synchronization accuracy. The greater vulnerability arises from the shared utilization of time and frequency resources by multiple users, making the received signal susceptible to the aggregate effects of CFOs from different user-dAP pairs. As a result, conventional CFO correction techniques—which typically assume a single dominant offset—become ineffective in distributed multi-user transmissions, where CFO contributions accumulate across spatially separated links.
@@ V-C3 Connection to the Analytical Framework
In summary, the experiments were conducted on the RENEW indoor massive MIMO testbed [31], using 64 SDRs grouped into four dAPs with 16 antennas each. The system operated in the CBRS band using OFDM with 64 subcarriers and $\Delta f=78.125$ kHz of spacing. We evaluated both single-user and multi-user uplink configurations under HUB-mode and LO-mode synchronization, with users remaining stationary during each experiment. For every scenario, 2000 transmission frames were recorded, and channel estimates and inter-dAP CFO values were extracted from the received pilot signals. The datasets are available at [32], and the corresponding codebase is available at[53]. These empirical values were incorporated into our simulation and analytical models for comparison. The results confirm that our CFO-aware SINR framework and simulation model provide accurate predictions of beamforming performance, particularly in capturing the degradation caused by inter-dAP CFO in distributed multi-user settings.
@@ V-D Discussion
The insights gained from analyzing the CFO profile, including its distribution and variance, can guide the design of key system parameters. These parameters might include subcarrier spacing, the number of beamforming antennas, and the selected beamforming approach, all of which are crucial for enabling multi-user beamforming with distributed antenna arrays. We have provided experimental datasets to evaluate the impact of CFO in real-world systems, available at [32] for further research. In our distributed scenario, while we refer to the placement of APs in Fig. 6a as distributed, they are only 3 meters apart. This setup may not fully capture the characteristics of a truly distributed system with diverse channel conditions. The main difference from a centralized case lies in the correlation pattern between antennas. However, our focus is not on the effects of varying channel conditions, diversity patterns, or scattering scenarios. Instead, we concentrate on the impact of CFO between distributed APs. In this context, the dAPs may be physically close or have correlated channels, but they operate with independent LOs. Thus, the word distributed in our study concentrates mainly on the use of independent RF frequency sources, potentially resulting in random frequency offsets between distributed nodes.
In our experiments in the LO-mode, we let every dAP use its own LO for deriving carrier and sampling frequencies, and there is no additional OTA frequency synchronization step performed. However, in practical systems, there is always a frequency synchronization mechanism to ensure phase coherency among radios participating in communication. We performed our experiments with free-running LOs to evaluate the performance in the worst-case scenario. An extra synchronization step can compensate for the CFO-induced residual phase offset for some duration of time, depending on the synchronization algorithm. To maintain phase coherence over time, the synchronization must be performed frequently, introducing communication overhead on the system that scales with the network size. Moreover, the CFO-induced residual phase offset accumulates in distributed MU-MIMO systems as the signal is affected by multiple CFO levels from all dAP participants. Several solutions and algorithms have been proposed in the literature for synchronization in distributed systems [55, 56, 7, 8], some of which have shown effectiveness in enabling D-MUBF on a small scale [17, 3, 4, 5]. Nevertheless, this work provides valuable insight into the impact of CFO on beamforming performance in distributed multi-user systems operating without perfect frequency synchronization.
@@ VI Conclusion
Appendix A Derivation of Frequency Doman OFDM Signal
@@
$\textstyle=\frac{1}{N_{sc}}\sum_{\zeta=0}^{N_{sc}-1}e^{j\frac{2\pi}{N_{sc}}(\zeta+n(N_{sc}+L_{cp}))\epsilon}$
@@
$\textstyle=e^{j\overline{\phi}_{m}[n]}\sum_{q=0}^{N_{sc}-1}\mathrm{s}_{k}[n,q]\mathrm{h}_{km}^{(i)}[q]G_{\epsilon}[q-l],$
@@
$\textstyle\quad+e^{\overline{\phi}_{m}[n]}\sum_{q\neq l}^{N_{sc}-1}\mathrm{s}[n,q]\mathrm{h}_{km}^{(i)}[q]G_{\epsilon}[q-l].$ (30)

### v2

@@ IV-B Monte Carlo Simulation of SINR Performance
We simulate an OFDM transmission and reception chain based on the 802.11 standard [802-11_standard] OFDM format. Out of $N_{sc}=64$ subcarriers, we use $48$ subcarriers for data transmission, $4$ for pilot, and the rest as null subcarriers. The model simulates $M=4$ dAPs with $N=16$ antennas per dAP, serving $K=4$ single-antenna users at the same time-frequency resource. Users transmit their signals in uplink with the same power $p_{\mathrm{ul}}$, where the ratio of $p_{\mathrm{ul}}$ to the noise variance $\sigma_{z}^{2}$ in the environment is $20$ dB. Users first send orthogonal uplink pilots. These pilots are composed of two consecutive long training sequence (LTS) symbols based on 802.11 format and will be used for channel estimation and combining weight calculation. The cyclic prefix length is $L_{cp}=16$, and QPSK modulation will be used, which is more robust to CFO compared to higher-order modulations. CFO is induced in the system in the form of realizations of the normal distribution as $\epsilon\sim\mathcal{N}(0,\beta^{2})$. After receiving signals and estimating channels, CBF is applied, and SINR and EVM are measured for each scenario.
@@
These simulations are conducted to compare the measured SINR, obtained by introducing CFO as random realizations, with the analytical SINR expression in (III-B2), where the CFO effects are modeled using the closed-form approximations in (17)–(22). A comparison is made in Fig. 3a, where both the measured SINR based on the Monte Carlo simulation (labeled as Sim) and the analytical SINR based on (III-B2) and (17)–(22) (labeled as Anlt) are shown for CBF. The CFO distribution is normal, with STD $\beta$ ranging from $0.5\%$ to $5\%$. For instance, when $\beta=2\%$, according to the characteristics of the normal distribution, over $99.7\%$ of the normalized CFO ($\epsilon$) values are within $3\beta=6\%$ of the mean, which means $\epsilon\in[-6\%,\,6\%]$. For a subcarrier spacing of $\Delta f=30$ kHz, $6\%$ normalized CFO corresponds to an offset of $\delta_{f}=0.06\times\Delta f=1.8$ kHz from the true carrier frequency. The ratio of this offset to the carrier frequency ($\delta_{f}/f_{c}$), measured in parts per billion (PPB), is equivalent to $750$ PPB, $600$ PPB, and $500$ PPB for the center frequencies of $2.4$ GHz, $3$ GHz, and $3.6$ GHz, respectively. This range of CFO can significantly affect the performance of the system. The results in Fig. 3a illustrate that SINR drops with increasing $\beta$, and this drop is faster for larger OFDM symbol indices. This is because the phase rotation caused by CFO linearly scales with the symbol index $n$ within an uplink slot with multiple OFDM symbols. It also demonstrates that the SINR expression (III-B2) with the CFO model (17)–(22) gives an appropriate estimate of the average SINR for different $\beta$ and $n$ values.
Although we assess the SINR in our analytical framework, EVM is often used as a performance metric to evaluate empirical datasets [nakagawa_phase_noise, helfenstein_evm]. SINR accounts for interference effects and serves to measure the channel quality indicator (CQI), which is utilized as the parameter for dynamically adjusting the modulation and coding scheme (MCS) [asilomar_mcs, bobrov_mcs]. However, in a multi-user transmission scenario, measuring the power of IUI and separating it from the power of a desired signal and the power of noise is not feasible in most practical scenarios, subsequently making accurate SINR measurements difficult [brown_evm, marco_channel_dynamics]. Due to the practical challenges of SINR measurement, EVM can be a viable alternative metric in real-world experiments because of the feasibility of measurement. EVM can provide insight into the different sources of imperfections, including carrier leakage, IQ imbalance, non-linearity, and LO frequency errors [hisham_evm, liu_evm_est]. Moreover, in several wireless communication standards, the requirements on EVM are defined as part of the standard, such as IEEE 802.11a [ieee_wlan] and IEEE 802.16e [ieee_wimax], among others.
@@
Accordingly, we evaluate the root mean square error vector magnitude ($\mathrm{EVM}_{rms}$) in our Monte Carlo simulations, as shown in Fig. 3b, where $\mathrm{EVM}_{rms}$ is defined as in [hisham_evm]:
@@ V Experiments
In this section, we first introduce the RENEW [renew] indoor massive MIMO platform. We then describe the experimental setup and outline the datasets collected under various experimental conditions. Furthermore, we showcase several experimental results derived from these datasets along with comparisons with the analytical framework discussed in § III.
@@ V-A Hardware Setup
We use the RENEW programmable massive MIMO platform [renew] to perform experiments. The RENEW base station uses the Iris SDR [shepard2020argosnet] as its building block, with each SDR supporting two RF chains. The base station includes a central hub that by default distributes clock and time trigger signals to up to 48 Iris SDRs (96 antennas) for time and frequency synchronization. At each Iris, both the RF and sampling clocks are derived from the clock distributed by the hub. Therefore, all antennas are locked at the carrier frequency $f_{c}$ in the form of a perfectly synchronized array. In this scenario, perfect time and frequency synchronization among antennas is achieved via the central hub. We denote this operational mode of the SDRs as HUB-mode. An Iris can also be used in a standalone mode to emulate a user. The RENEW platform is designed to operate at the CBRS spectrum ranging from 3.553.55-$3.7$ GHz.
Apart from receiving an external clock, each Iris is equipped with its own LO, a super-TCXO with a nominal $\pm 100$ PPB accuracy for the sub-$6$ GHz band. Through firmware settings, the Iris can be configured to use its own LO to generate the RF clock or to use the external clock from the central hub. To emulate a distributed antenna array with independent LOs, each Iris in the base station can be programmed to use its own LO, introducing a random CFO between the antennas. In this operation mode, the antennas still receive the time trigger signal from the central hub as their time reference. Time synchronization in distributed systems can be achieved through well-established protocols, such as IEEE 1588 Precision Time Protocol (PTP) or GPS-based methods [sami2010time_sync, schwartz2022time_sync, chen2019time_sync, zhang2023time_sync]. Distributing the time trigger across radios ensures time alignment, allowing us to isolate and focus solely on the impact of CFO on system performance. We denote the operation mode in which SDRs use their LO to derive the RF clock as LO-mode.
@@ V-B Experiment Design
We use the RENEWLab real-time software [RENEWLab] to collect datasets in all of the configurations. The RENEWLab software runs on a host server that is connected to the RENEW BS through fiber. Through a configurable frame schedule, RENEWLab software orchestrates the transmission and reception of pilot and data OFDM symbols at users and from the BS antennas, respectively. Within each frame, one of the BS antennas transmits a synchronization beacon so that users can synchronize their frame time boundaries with the BS frame. We set up the frame schedule for the users so that each user sends time-orthogonal pilots and then all simultaneously transmit data OFDM symbols. We use the 802.11 LTS signal as the pilot symbol, which is used for channel estimation and beamforming matrix calculation. The format of the data OFDM symbols also follows the 802.11 standard [802-11_standard], where the FFT size is $64$. Out of the $64$ subcarriers, $48$ subcarriers carry data, $4$ are pilot subcarriers, and the remaining are null subcarriers. The pilot subcarriers are used for residual phase offset correction. We try QPSK for modulating data subcarriers. The detailed parameters of the experiments are shown in Table I. The RENEWLab framework enables the recording of pilot and data OFDM symbols received at BS antennas in a dataset file with HDF5 format [hdf5]. The HDF5 datasets can be processed offline with the Python-based post-processing library in the RENEWLab software. Our data collection effort resulted in over 250 HDF5 files, with each file comprising 2000 frames, and each frame containing pilot symbols for each user followed by $10$ data OFDM symbols. Datasets are available on [dist_mu_mimo_datasets] for further research purposes.
@@ V-C2 Beamforming Performance Analysis
Across various multi-user configurations, where different user subsets were selected to capture the effects of high and low spatial correlation in beamforming, we measured the average estimated CFO between the dAPs. The results show that inter-dAP CFO values in these multi-user experiments ranged from $150$ to $350$ PPB, while intra-dAP synchronization remained consistently maintained, as expected. Notably, the maximum observed CFO across all multi-user measurements approached approximately $900$ PPB, which corresponds to a normalized CFO of $\epsilon=4.15\,\%$ when considering a subcarrier spacing of $\Delta f=78.125$ kHz. These results are consistent with our assumption of negligible ICI impact, given that normalized inter-dAP CFO variations remain within the $5\%$ range.
@@
For two multi-user configurations of $64\times 2$ and $64\times 4$ with distributed placement of APs, we processed the received data samples in both the HUB-mode and LO-mode, employing ZFBF to mitigate inter-user interference and recover the uplink data from each user. In the corresponding Monte Carlo simulations, we used the estimated CSI and CFO values obtained from the experimental measurements. The average per-user EVM-SNR measurements for these multi-user scenarios are presented in Fig. 9, comparing the performance of Dist-HUB versus Dist-LO scenarios. As illustrated in Fig. 9, the performance drop between the Dist-HUB and Dist-LO scenarios is notable. In the two-user configuration, there is a reduction of approximately $5$ dB in the average EVM-SNR, while this degradation escalates to approximately $9$ dB in the four-user setup. This performance gap highlights the increased sensitivity of higher-order multi-user MIMO configurations to synchronization accuracy. The greater vulnerability arises from the shared utilization of time and frequency resources by multiple users, making the received signal susceptible to the aggregate effects of CFOs from different user-dAP pairs. As a result, conventional CFO correction techniques, which typically assume a single dominant offset, become ineffective in distributed multi-user transmissions, where CFO contributions accumulate across spatially separated links.
@@ V-C3 Connection to the Analytical Framework
In summary, the experiments were conducted on the RENEW indoor massive MIMO testbed [renew], using 64 SDRs grouped into four dAPs with 16 antennas each. The system operated in the CBRS band using OFDM with 64 subcarriers and $\Delta f=78.125$ kHz of spacing. We evaluated both single-user and multi-user uplink configurations under HUB-mode and LO-mode synchronization, with users remaining stationary during each experiment. For every scenario, 2000 transmission frames were recorded, and channel estimates and inter-dAP CFO values were extracted from the received pilot signals. The datasets are available at [dist_mu_mimo_datasets], and the corresponding codebase is available at[RENEWLab]. These empirical values were incorporated into our simulation and analytical models for comparison. The results confirm that our CFO-aware SINR framework and simulation model provide accurate predictions of beamforming performance, particularly in capturing the degradation caused by inter-dAP CFO in distributed multi-user settings.
@@ V-D Discussion
The insights gained from analyzing the CFO profile, including its distribution and variance, can guide the design of key system parameters. These parameters might include subcarrier spacing, the number of beamforming antennas, and the selected beamforming approach, all of which are crucial for enabling multi-user beamforming with distributed antenna arrays. We have provided experimental datasets to evaluate the impact of CFO in real-world systems, available at [dist_mu_mimo_datasets] for further research. In our distributed scenario, while we refer to the placement of APs in Fig. 6a as distributed, they are only 3 meters apart. This setup may not fully capture the characteristics of a truly distributed system with diverse channel conditions. The main difference from a centralized case lies in the correlation pattern between antennas. However, our focus is not on the effects of varying channel conditions, diversity patterns, or scattering scenarios. Instead, we concentrate on the impact of CFO between distributed APs. In this context, the dAPs may be physically close or have correlated channels, but they operate with independent LOs. Thus, the word distributed in our study concentrates mainly on the use of independent RF frequency sources, potentially resulting in random frequency offsets between distributed nodes.
In our experiments in the LO-mode, we let every dAP use its own LO for deriving carrier and sampling frequencies, and there is no additional OTA frequency synchronization step performed. However, in practical systems, there is always a frequency synchronization mechanism to ensure phase coherency among radios participating in communication. We performed our experiments with free-running LOs to evaluate the performance in the worst-case scenario. An extra synchronization step can compensate for the CFO-induced residual phase offset for some duration of time, depending on the synchronization algorithm. To maintain phase coherence over time, the synchronization must be performed frequently, introducing communication overhead on the system that scales with the network size. Moreover, the CFO-induced residual phase offset accumulates in distributed MU-MIMO systems as the signal is affected by multiple CFO levels from all dAP participants. Several solutions and algorithms have been proposed in the literature for synchronization in distributed systems [gao2019_syncAlg, wang_2015_syncAlg, ahmad_2024_syncAlg, milos_2023_syncAlg], some of which have shown effectiveness in enabling D-MUBF on a small scale [chorus, rogalin, rashid_2022_syncAlg, Rashid_2023_syncAlg]. Nevertheless, this work provides valuable insight into the impact of CFO on beamforming performance in distributed multi-user systems operating without perfect frequency synchronization.
@@ VI Conclusion
Appendix A Derivation of Frequency Domain OFDM Signal
@@
$\textstyle=\frac{1}{N_{sc}}\sum_{\zeta=0}^{N_{sc}-1}e^{j\frac{2\pi}{N_{sc}}(\zeta+n(N_{sc}+L_{cp}))\epsilon_{m}}\cdot$
$\textstyle\quad\sum_{q=0}^{N_{sc}-1}\mathrm{s}_{k}[n,q]\mathrm{h}_{km}^{(i)}[q]e^{j\frac{2\pi\zeta}{N_{sc}}(q-l)}$
@@
$\textstyle=e^{j\overline{\phi}_{m}[n]}\sum_{q=0}^{N_{sc}-1}\mathrm{s}_{k}[n,q]\mathrm{h}_{km}^{(i)}[q]G_{\epsilon}[q-l]$
@@
$\textstyle\quad+e^{j\overline{\phi}_{m}[n]}\sum_{q\neq l}^{N_{sc}-1}\mathrm{s}[n,q]\mathrm{h}_{km}^{(i)}[q]G_{\epsilon}[q-l].$ (30)